foxey3:

how many radians are in 1300

Luigi0210:

One radian is defined as \(\large \frac{\pi}{180} \) and by 1300, I assume you mean \( 1300°\) so : \(\Large 1300* \frac{\pi}{180} = \frac{65}{9}*\pi \)